Formulate a system of equations for the situation below and solve.
A theater has a seating capacity of 900 and charges $4 for children, $6 for students, and $8 for adults. At a certain screening with full attendance, there were half as many adults as children and students combined. The receipts totaled $5200. How many children attended the show?

Answer by ikleyn(52814)

.
Formulate a system of equations for the situation below and solve.
A theater has a seating capacity of 900 and charges $4 for children, $6 for students, and $8 for adults.
At a certain screening with full attendance, there were half as many adults as children and students combined.
The receipts totaled $5200. How many children attended the show?
~~~~~~~~~~~~~~~~~~~~~~~~

Since "there were half as many adults as children and students combined", you can momentarily conclude that there were 300 adults 
(1/3 of the full capacity) and 600 children and students combined (2/3 of the full capacity).


Let C be the number of children.

Then the number of students was  (600-C).


The "value" equation for the receipts totaled is THIS:

4C + 6*(600-C) + 8*300 = 5200      dollars.


    *** One single equation for one single unknown. ***


4C + 3600 - 6C + 2400 = 5200

-2C = 5200 - 2400 - 3600 = -800  ====>  C = 400.


Answer.  400 children.

----------------
Notice.  This problem is for one single equation,  actually.

Therefore,  my major wish was do not make it more complicated than it really should be.